OpenStudy (anonymous):

solve. sqrt 5^8n=125^(n+4)

Parth (parthkohli):

Get the same base first of all. \( \color{Black}{\Rightarrow 5^{8n}= (5^3)^{n + 4} }\) Now you can equate the exponents. \( \color{Black}{\Rightarrow 8n = 3(n + 4) }\)

OpenStudy (anonymous):

yup @ParthKohli is correct!

OpenStudy (anonymous):

8n = 3n +12 5n=12 n=12/5

OpenStudy (anonymous):

@Jshakour give medal to @ParthKohli he deserves it!

OpenStudy (anonymous):

okay so you can just get rid of the sqrt sign in the beginning?

OpenStudy (anonymous):

Its \[\sqrt{5} \]

OpenStudy (anonymous):

@Jshakour u only need to find the value of n!!!

OpenStudy (anonymous):

but in the practice test 12/5 isn't an option...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

the options are 12, 4 and -4 and -12

OpenStudy (anonymous):

but @Jshakour that cant be correct